# Long Chain-of-Thought for Reasoning Large Language Models

### v3

@@ 8.2 Multilingual Long CoT
While significant progress has been made in RLLMs for the English language, expanding reasoning capabilities to multiple languages is essential for the creation of RLLMs that can effectively perform complex, multi-step tasks across a variety of linguistic contexts [493, 494, 162]. Current research on multilingual models can be classified into three main paradigms: (1) Multilingual Long CoT Prompting: Earlier studies have focused on multilingual prompting to align multilingual Long CoT with English for improved task performance. For instance, XLT [219] and CLP [490] employ generic template prompts that stimulate both cross-lingual and logical reasoning skills, enhancing task performance across languages. (2) Multilingual Long CoT Training: Researchers have proposed multilingual SFT or RL methods to improve reasoning consistency across languages. Notable examples include the mCoT [346] and xCoT [57] frameworks, which align reasoning processes between high- and low-resource languages. Additionally, the DRT-o1 [622] method extends the success of Long CoT to neural machine translation. More recently, Wang et al. [642] suggest that training multilingual PRMs on diverse datasets can enhance multi-step reasoning capabilities across linguistic backgrounds. (3) Multilingual Long CoT Test Time Scaling: Earlier, Qin et al. [490] first introduced CLSP as a method to scale reasoning tasks across different language speakers. Building on this foundation, AutoCAP [850] utilizes RLLMs as verifiers to automatically select languages and assign appropriate weights, facilitating a more diverse scaling approach. Furthermore, Ranaldi et al. [503] propose a tree search method to further enhance the depth of scaling.
@@ 8.3 Agentic & Embodied Long CoT
Researchers have expanded Long CoT in interactive environments, significantly improving success rates in automated exploration tasks for agentic systems [185, 869, 815, 635, 637, 370] and embodied systems [803, 700, 844]. Current research primarily focuses on two approaches: (1) Tree-based Search Augmentation Early work by Hao et al. [185], Koh et al. [283] introduce tree search techniques to enhance agent exploration. Hu et al. [212] further propose planning sampling strategies to accelerate tree search processes. Additionally, Light et al. [356] develop a method to gather high-quality interactive feedback through self-play simulations with MCTS and LLM-based reflection, which helps acquire high-level strategic skills and guide low-level execution. (2) Environmental Interactivity Improvement A key feature of Agentic Systems is their understanding for the physical world [24, 278] and interaction with the environment [881, 141], making the enhancement of this aspect a critical focus [185]. Nie et al. [447] and Hu et al. [211] improve interactivity by incorporating memory history into the agent’s functions. (3) Multi-agent Cooperative Improvement Another key feature of agentic systems is that it can incorporate multiple agents to cooperative to solve a complex problem [903, 625, 482, 690, 901, 606, 763]. Christakopoulou et al. [108] introduce the Talker-Reasoner architecture, which separates the agent’s tasks into deep reasoning and rapid dialogue generation, providing a more effective interaction protocol. Lei et al. [305] introduce the Multi-Agent System for Conditional Mining (MACM) prompting method, which effectively addresses complex mathematical problems and exhibits robust generalization across diverse mathematical contexts.
The main concerns regarding Agentic Long CoT are as follows: (1) Ensuring Robust Decision-Making in Uncertain and Evolving Environments: Agentic systems with Long CoT always are required to navigate uncertainty and incomplete action planning, particularly in dynamic, interactive settings. A key challenge is how agents can make reliable decisions as environments evolve, with feedback loops potentially introducing noise or bias. (2) Scalability and Efficiency Across Multi-Agent Interactions: A major concern is how agentic systems can scale multi-agent and reasoning processes in complex, long-term interactions. As agents engage in extended tasks, maintaining interaction efficiency while managing large volumes of data—such as memory history and real-time feedback—becomes increasingly difficult [38].
@@ 8.4 Efficient Long CoT
The deep reasoning, exploration, and reflection of the Long CoT often lead to long outputs, which necessitate improved speedup techniques [749, 218, 156, 542, 244, 311]. Consequently, optimizing reasoning for faster reasoning with maximum accuracy has become a significant challenge for Long CoT [157]. Current research mainly focuses on two approaches: (1) Direct Compression and Shortening of Reasoning Chains: The most direct strategy is to consider direct compression and reducing the length of the reasoning chain while maintaining accuracy [104, 550, 22, 205]. Specifically, a series of work [572, 406, 59, 417, 444] encourage the generation of shorter reasoning processes, minimizing redundancy and enhancing efficiency [19, 720, 31, 442, 639, 393]. Additionally, researchers further introduce token budgets in prompts to control reasoning complexity, further improving efficiency [183, 808, 607, 243, 319, 5]. Building on these approaches, MARP [78] and DynaThink [454] allow LLMs to adapt reasoning speed based on task complexity, perplexity, or confidence, optimizing both efficiency and accuracy [171, 521, 906, 122, 116, 632, 269]. Moreover, Botta et al. [48] and Xia et al. [694] introduce a technique that enables LLMs to erase or skip some generated tokens, thereby compressing the reasoning length. More radically, Yu et al. [780] and Du et al. [130] propose distilling long reasoning paradigms into direct prediction models, reducing computational costs without sacrificing reasoning quality. (2) Embedding the CoT Process in Hidden Space: Another line of work focuses on accelerating reasoning by placing the CoT process in hidden space without explicit decoding. Specifically, Coconut [187], LaTRO [66], and SoftCoT [723] transfer reasoning into continuous latent space, promoting "continuous thinking" and enabling the model to maintain multiple alternative reasoning paths [830]. Similarly, Wang et al. [646] use "planning tokens" to enhance reasoning, performing the planning process in hidden space to save computational resources and improve inference performance.
@@
The main concerns regarding efficiency for Long CoT are as follows: (1) Incorporating More Adaptive Reasoning Strategies: Future research should explore adaptive reasoning techniques that enable models to dynamically adjust the depth and complexity of Long CoT based on real-time evaluations of task difficulty and intermediate result quality [78, 352, 546, 791, 730, 529, 638], rather than relying solely on human experience. (2) Leveraging efficient reasoning format: Another promising direction involves integrating multimodal, latent space, or other efficient reasoning formats to express logic more effectively [102, 528]. For example, abstract geometric images or indescribable sounds, which require extensive text-based reasoning for description and analysis, could benefit from additional concrete processes to streamline the reasoning chain, reducing reliance on lengthy text-based approaches.
@@ 8.5 Knowledge-Augmented Long CoT
The reasoning model significantly enhances reasoning capabilities, but it still lacks knowledge in specialized fields and timely new information [80, 138, 379, 537]. Thus, enriching reasoning with additional knowledge presents a key challenge for Long CoT [71, 64]. Current research focuses primarily on two approaches: (1) Retrieval-Augmented Generation: Retrieval-Augmented Generation (RAG) techniques enhance LLMs by integrating dynamic knowledge retrieval and document refinement [337, 647, 174, 254, 661]. Research has combined RAG with reasoning modules to improve performance on complex tasks [576, 259, 378, 685, 76]. O1 Embedder [727] optimizes multi-task retrieval and reasoning through synthetic data training. Furthermore, Stream of Search (SoS) [149] CoRAG [631] boost search accuracy and addresses unresolved issues by incorporating more natural reflection and exploration in RAG. (2) Model Knowledge Injection: An alternative approach involves integrating additional knowledge during SFT or RL [390]. Specifically, HuatuoGPT-o1 [71] utilize the R1-like paradigm to train LLMs by model-judged reward RL, which significantly improves the medical knowledge during reasoning [457, 229]. Huang et al. [233] and Wang et al. [615] optimize for injecting medical knowledge in Long CoT scenarios by SFT, which also achieve great performance. Further, Jiang et al. [256] introduce MCTS to synthesize data, achieving superior performance. This model merges verifiable medical knowledge with reinforcement learning techniques to enhance performance in complex, medical task settings.
@@ 8.6 Safety and Stability for Long CoT
Despite the notable performance improvements brought about by Long CoT, Long CoT-augmented LLMs still encounter significant safety challenges. These include issues such as the generation of unstable outputs, exemplified by the tendency to memorize in-domain math questions instead of engaging in actual reasoning [726], and the production of unsafe outputs, such as misinformation and offensive content [890, 309, 889, 402, 20, 40, 39, 127, 275, 842]. Current research primarily addresses two key approaches: (1) Long CoT Attack Several studies show that Long CoT makes models more vulnerable to unexpected behavior [140] or unsafe outputs [288, 904, 719, 92, 18]. For instance, Arrieta et al. [21] identify that DeepSeek-R1 is prone to generating harmful content, including misinformation and offensive speech. Additionally, Kumar et al. [285] introduce the OverThink attack, which exploits false inference problems to induce overthinking in models, providing insights into potential defensive strategies. Further, Yao et al. [758] fool RLLMs chain of iterative chaos, for better jailbreaking. (2) Long CoT Safety Improvement Another major area of research focuses on enhancing safety [252, 900, 388] and reliability [565, 506, 598, 117, 89] through prompting [147] or training [459] techniques. Shen et al. [528] present Heima, which optimizes inference efficiency and robustness. Gallego [147] proposes dynamic security prompts during inference, while Cheng et al. [101] address hallucinations by guiding reasoning with a tree search algorithm. Zhao et al. [865] introduce a self-reflection framework to identify biases, and Wang et al. [620] propose Safety Reasoning with Guidelines (SRG) to defend against out-of-distribution attacks. Finally, Parmar and Govindarajulu [466] combine reinforcement learning (RL) and supervised fine-tuning (SFT) in a hybrid training approach to reduce harmful outputs and enhance DeepSeek-R1’s safety.
The main concerns regarding safety for Long CoT are as follows: (1) Mitigating Cognitive Overload in Complex Reasoning: Long CoT approaches require managing extended reasoning chains, which can result in cognitive overload in LLMs [260, 78]. This overload may lead to errors, hallucinations, or unsafe outputs. Developing strategies that allow LLMs to maintain accuracy and coherence during complex reasoning, without overwhelming their capacity, remains a key challenge for ensuring safety. (2) Balancing Model Performance with Safety: A major challenge lies in balancing improved model performance with safety [227]. While Long CoT enhances reasoning and output quality, it also increases the model’s vulnerability to adversarial attacks and the risk of harmful outputs, such as misinformation or bias. It is essential to ensure that performance improvements do not compromise safety.
@@ 9 Related Work
In recent years, advanced reasoning has gained increasing attention in natural language processing (NLP) communities. Early works, such as Plaat et al. [478], Huang and Chang [222], and Chu et al. [110], explore the emergence of reasoning abilities in RLLMs as they scale, focusing on their capacity for in-context and few-shot learning across a range of tasks. Additionally, Giadikiaroglou et al. [163], Yu et al. [777] and Liu et al. [377] provide comprehensive overviews of LLM advancements in various reasoning tasks [549]. Moreover, Chu-Carroll et al. [111] highlight the need for hybrid architectures to address LLMs’ reliance on statistical patterns over structured reasoning.
With the development of advanced RLLMs, such as OpenAI-O1 and DeepSeek-R1, recent research has focused on improving reasoning capabilities. Patil [467] highlight the limitations of standard LLMs in addressing complex reasoning tasks, such as optimization and multi-step reasoning. In addition, Liang et al. [351] and Li [338] review strategies to scale search and test time, including the use of algorithms like Monte Carlo Tree Search, to enhance LLM reasoning. Xu et al. [712] examine the role of reinforcement learning and "thought" sequences in reasoning improvement [287], while Hong et al. [201] demonstrate the impact of prompting techniques. Further, Liu et al. [377] and Mondorf and Plank [439] stress the importance of deeper analysis beyond surface-level accuracy, and He et al. [195] explore self-evolutionary processes as a means to advance LLM reasoning. Besta et al. [44] propose a modular framework integrating structure, strategy, and training methods as part of a comprehensive system design approach. Most recently, Li et al. [347] provide a systematic survey of System 2 thinking, focusing on the methods used to differentiate them from System 1 thinking.

### v4

@@ 8.2 Multilingual Long CoT
While significant progress has been made in RLLMs for the English language, expanding reasoning capabilities to multiple languages is essential for the creation of RLLMs that can effectively perform complex, multi-step tasks across a variety of linguistic contexts [594, 596, 194, 66, 760]. Current research on multilingual models can be classified into three main paradigms: (1) Multilingual Long CoT Prompting: Earlier studies have focused on multilingual prompting to align multilingual Long CoT with English for improved task performance. For instance, XLT [264] and CLP [591] employ generic template prompts that stimulate both cross-lingual and logical reasoning skills, enhancing task performance across languages. (2) Multilingual Long CoT Training: Researchers have proposed multilingual SFT or RL methods to improve reasoning consistency across languages [747]. Notable examples include the mCoT [410] and xCoT [62] frameworks, which align reasoning processes between high- and low-resource languages. Additionally, the DRT-o1 [746] method extends the success of Long CoT to neural machine translation. More recently, Wang et al. [774] suggest that training multilingual PRMs on diverse datasets can enhance multi-step reasoning capabilities across linguistic backgrounds. (3) Multilingual Long CoT Inference-Time Scaling: Earlier, Qin et al. [591] first introduced CLSP as a method to scale reasoning tasks across different language speakers. Building on this foundation, AutoCAP [1030] utilizes RLLMs as verifiers to automatically select languages and assign appropriate weights, facilitating a more diverse scaling approach. Furthermore, Ranaldi et al. [607] propose a tree search method to further enhance the depth of scaling.
@@ 8.3 Agentic & Embodied Long CoT
Researchers have expanded Long CoT in interactive environments by utilizing tools, significantly improving success rates in automated exploration tasks [220, 1057, 985, 167, 575]. Current research primarily focuses on two approaches: (1) Tree-based Search Augmentation Early work [220, 336] introduce tree search techniques to enhance agent exploration. Hu et al. [253] further propose planning sampling strategies to accelerate tree search processes. Additionally, Light et al. [424] develop a method to gather high-quality interactive feedback through self-play simulations with MCTS and LLM-based reflection, which helps acquire high-level strategic skills and guide low-level execution. (2) Environmental Interactivity Improvement A key feature of Agentic Systems is their understanding for the physical world [26, 331] and interaction with the environment [1072, 170, 640, 457], making the enhancement of this aspect a critical focus [220, 1072, 331, 170]. Nie et al. [540] and Hu et al. [252] improve interactivity by incorporating memory history into the agent’s functions. (3) Multi-agent Cooperative Improvement Another key feature of agentic systems is that it can incorporate multiple agents to cooperative to solve a complex problem [1100, 750, 581, 837, 1097, 729, 928]. Christakopoulou et al. [130] introduce the Talker-Reasoner architecture, which separates the agent’s tasks into deep reasoning and rapid dialogue generation, providing a more effective interaction protocol. Lei et al. [360] introduce the Multi-Agent System for Conditional Mining (MACM) prompting method, which effectively addresses complex mathematical problems and exhibits robust generalization across diverse mathematical contexts.
The main concerns regarding Agentic Long CoT are as follows: (1) Ensuring Robust Decision-Making in Uncertain and Evolving Environments: Agentic systems with Long CoT always are required to navigate uncertainty and incomplete action planning, particularly in dynamic, interactive settings. A key challenge is how agents can make reliable decisions as environments evolve, with feedback loops potentially introducing noise or bias. (2) Scalability and Efficiency Across Multi-Agent Interactions: A major concern is how agentic systems can scale multi-agent and reasoning processes in complex, long-term interactions [256]. As agents engage in extended tasks, maintaining interaction efficiency while managing large volumes of data—such as memory history and real-time feedback—becomes increasingly difficult [40, 945].
@@ 8.4 Efficient Long CoT
The deep reasoning, exploration, and reflection of the Long CoT often lead to long outputs, which necessitate improved speedup techniques [188, 658, 471, 600, 168, 469, 638, 794], such as KV Cache optimization [999, 911, 464], token compression [507, 537, 960, 200, 875, 163, 651], efficient structure [293, 263, 114, 65, 235, 354, 554, 877] and dynamic reasoning patterns [758, 145, 664, 480, 367, 307, 1017, 826, 436, 449, 846, 329, 935, 719, 1023]. Consequently, optimizing reasoning for faster reasoning with maximum accuracy has become a significant challenge for Long CoT [189, 1046]. Current research mainly focuses on two approaches: (1) Direct Compression and Shortening of Reasoning Chains: The most direct strategy is to consider direct compression and reducing the length of the reasoning chain while maintaining accuracy [124, 669, 24, 246, 541, 940, 467, 117]. Specifically, a series of work [694, 493, 64, 507, 537] encourage the generation of shorter reasoning processes [33, 535, 771, 186] or removing reflection signal tokens [735], minimizing redundancy and enhancing efficiency [21, 873, 476]. Additionally, researchers further introduce token budgets in prompts to control reasoning complexity, further improving efficiency [218, 978, 730, 292, 375, 5, 408]. Building on these approaches, MARP [85] and DynaThink [548] allow LLMs to adapt reasoning speed based on task complexity, perplexity, or confidence, optimizing both efficiency and accuracy [204, 627, 1105, 145, 138, 758, 321, 465, 313, 832, 1101]. Moreover, Botta et al. [51] and Xia et al. [842] introduce a technique that enables LLMs to erase or skip some generated tokens, thereby compressing the reasoning length [1103]. More radically, Yu et al. [947] and Du et al. [153] propose distilling long reasoning paradigms into direct prediction models, reducing computational costs without sacrificing reasoning quality. (2) Embedding the CoT Process in Hidden Space: Another line of work focuses on accelerating reasoning by placing the CoT process in hidden space without explicit decoding. Specifically, Coconut [222], LaTRO [73], and SoftCoT [879] transfer reasoning into continuous latent space, promoting “continuous thinking” and enabling the model to maintain multiple alternative reasoning paths [1003, 880]. Similarly, Wang et al. [780] use "planning tokens" to enhance reasoning, performing the planning process in hidden space to save computational resources and improve inference performance.
@@
The main concerns regarding efficiency for Long CoT are as follows: (1) Incorporating More Adaptive Reasoning Strategies: Future research should explore adaptive reasoning techniques that enable models to dynamically adjust the depth and complexity of Long CoT based on real-time evaluations of task difficulty and intermediate result quality [85, 419, 663, 959, 888, 636, 769, 273, 761] or even diffusion-like decoding processes [344], rather than relying solely on human experience. (2) Leveraging efficient reasoning format: Another promising direction involves integrating multimodal, latent space, or other efficient reasoning formats to express logic more effectively [120, 635, 770]. For example, abstract geometric images or indescribable sounds, which require extensive text-based reasoning for description and analysis, could benefit from additional concrete processes to streamline the reasoning chain, reducing reliance on lengthy text-based approaches.
@@ 8.5 Knowledge-Augmented Long CoT
The reasoning model significantly enhances reasoning capabilities, but it still lacks knowledge in specialized fields and timely new information [88, 165, 452, 650]. Thus, enriching reasoning with additional knowledge presents a key challenge for Long CoT [78, 71]. Current research focuses primarily on two approaches: (1) Retrieval-Augmented Generation: Retrieval-Augmented Generation (RAG) techniques enhance LLMs by integrating dynamic knowledge retrieval and document refinement [397, 781, 207, 303, 797, 1061, 1058, 566]. Research has combined RAG with reasoning modules to improve performance on complex tasks [698, 310, 451, 828, 83, 1020, 590]. O1 Embedder [884] optimizes multi-task retrieval and reasoning through synthetic data training. Furthermore, Stream of Search (SoS) [180], and CoRAG [757] boost search accuracy and addresses unresolved issues by incorporating more natural reflection and exploration in RAG. (2) Model Knowledge Injection: An alternative approach involves integrating additional knowledge during SFT or RL [473, 993, 119, 1090]. Specifically, HuatuoGPT-o1 [78] utilize the R1-like paradigm to train LLMs by model-judged reward RL, which significantly improves the medical knowledge during reasoning [551, 277]. Huang et al. [282] and Wang et al. [739] optimize for injecting medical knowledge in Long CoT scenarios by SFT, which also achieve great performance. Further, Jiang et al. [306] introduce MCTS to synthesize data, achieving superior performance. This model merges verifiable medical knowledge with reinforcement learning techniques to enhance performance in complex, medical task settings.
@@ 8.6 Safety and Stability for Long CoT
Despite the notable performance improvements brought about by Long CoT, Long CoT-augmented LLMs still encounter significant safety and stability challenges [1093, 1032, 492, 807]. These include issues such as the generation of unstable outputs, exemplified by the tendency to memorize in-domain math questions instead of engaging in actual reasoning [883], and the production of unsafe outputs, such as misinformation and offensive content [1081, 365, 1080, 487, 22, 42, 41, 150, 327, 1021]. Current research primarily addresses two key approaches: (1) Long CoT Attack Several studies show that Long CoT makes models more vulnerable to unexpected behavior [169, 139], hallucinations [239, 482] or unsafe outputs [341, 1102, 872, 103, 19, 502]. For instance, Arrieta et al. [23] identify that DeepSeek-R1 is prone to generating harmful content, including misinformation and offensive speech. Additionally, Kumar et al. [338] introduce the OverThink attack, which exploits false inference problems to induce overthinking in models, providing insights into potential defensive strategies. Further, Yao et al. [922] fool RLLMs chain of iterative chaos, for better jailbreaking. (2) Long CoT Safety Improvement Another major area of research focuses on enhancing safety [301, 1095, 470] and reliability [687, 610, 721, 140, 100, 628] through prompting [178] or training [553] techniques. Shen et al. [635] present Heima, which optimizes inference efficiency and robustness. Gallego [178] proposes dynamic security prompts during inference, while Cheng et al. [116] address hallucinations by guiding reasoning with a tree search algorithm. Zhao et al. [1050] introduce a self-reflection framework to identify biases, and Wang et al. [744] propose Safety Reasoning with Guidelines (SRG) to defend against out-of-distribution attacks. Finally, Parmar and Govindarajulu [561] combine reinforcement learning (RL) and supervised fine-tuning (SFT) in a hybrid training approach to reduce harmful outputs and enhance DeepSeek-R1’s safety.
The main concerns regarding safety for Long CoT are as follows: (1) Mitigating Cognitive Overload in Complex Reasoning: Long CoT approaches require managing extended reasoning chains, which can result in cognitive overload in LLMs [311, 85]. This overload may lead to errors, hallucinations, or unsafe outputs. Developing strategies that allow LLMs to maintain accuracy and coherence during complex reasoning, without overwhelming their capacity, remains a key challenge for ensuring safety and trustworthiness [112]. (2) Balancing Model Performance with Safety: A major challenge lies in balancing improved model performance with safety [275]. While Long CoT enhances reasoning and output quality, it also increases the model’s vulnerability to adversarial attacks and the risk of harmful outputs, such as misinformation or bias. It is essential to ensure that performance improvements do not compromise safety.
@@ 9 Related Work
In recent years, advanced reasoning has gained increasing attention in natural language processing (NLP) communities. Early works [577, 268, 132], explore the emergence of reasoning abilities in RLLMs as they scale, focusing on their capacity for in-context and few-shot learning across a range of tasks. Additionally, Giadikiaroglou et al. [195], Yu et al. [943] and Liu et al. [450] provide comprehensive overviews of LLM advancements in various reasoning tasks [668]. Moreover, Chu-Carroll et al. [133] highlight the need for hybrid architectures to address LLMs’ reliance on statistical patterns over structured reasoning.
With the development of advanced RLLMs, such as OpenAI-o1 and DeepSeek-R1, recent research has focused on improving reasoning capabilities, especially on mathematical reasoning [765, 1054]. Patil [562] highlight the limitations of standard LLMs in addressing complex reasoning tasks, such as optimization and multi-step reasoning. In addition, Liang et al. [417] and Li [398] review strategies to scale search and inference time, including the use of algorithms like Monte Carlo Tree Search, to enhance LLM reasoning. Xu et al. [865] examine the role of reinforcement learning and "thought" sequences in reasoning improvement [340], while Hong et al. [242] demonstrate the impact of prompting techniques. Further, Liu et al. [450] and Mondorf and Plank [532] stress the importance of deeper analysis beyond surface-level accuracy, and He et al. [233] explore self-evolutionary processes as a means to advance LLM reasoning. Besta et al. [46] propose a modular framework integrating structure, strategy, and training methods as part of a comprehensive system design approach. Most recently, Li et al. [411] provide a systematic survey of System 2 thinking, focusing on the methods used to differentiate them from System 1 thinking.